OpenStudy (anonymous):

FUN TIME WITH MATH >:[ DUUUUU ITTTT! NOW! Three men walk inside a hotel and decide to stay for the night. They pay the price of $30 by splitting it so they each pay $10. They go up their room, but then the manager realizes that the price that night is $25 dollars so he sends $5 dollars up with the bell boy. The bellboy realizes that $5 can't be split by 3 so he pockets $2 and gives the men $1 each. But if you calculate it, the 3 men each payed $9, and 9 times 3 is 27, then add the $2 from the bellboy you got $29. Where did the last $1 go?

OpenStudy (asnaseer):

$27 already includes the $2 that the bellboy kept. you need to add the $3 they had left to $27 to get $30.

OpenStudy (anonymous):

asnaseer i need your help

OpenStudy (anonymous):

awee :/ -looka for a hardre one-

OpenStudy (anonymous):

looks for a harder one*

OpenStudy (asnaseer):

:-) @moneybird - how can I help?

OpenStudy (anonymous):

proving

OpenStudy (lilg132):

9 * 3 = 27 plus the 3 pounds they all recieved which gives you 30 or 9 * 3 = 27 - 2 pounds the bellboy took gives you 25 pounds